# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ Appendix B Inhomogeneous Maurer-Cartan equation
$\partial{\mathsf{q}}=\frac{1}{2}[{\mathsf{p}}_{2},{\mathsf{p}}_{2}]-\frac{1}{2}[{\mathsf{p}}_{1},{\mathsf{p}}_{1}]=[{\mathsf{q}},{\mathsf{p}}_{1}]+\frac{1}{2}[{\mathsf{q}},{\mathsf{q}}]=0.$
By acyclicity of ${\mathfrak{j}}_{N}$, we have ${\mathsf{q}}=\partial{\mathsf{a}}$ for some ${\mathsf{a}}\in{\mathfrak{j}}_{N,0}$. Then Lemma B.5 implies that $\exp({\mathsf{a}})\in\exp({\mathfrak{M}}_{N,0})$ maps ${\mathsf{p}}_{2}$ to ${\mathsf{p}}_{1}$. ∎
Lemma B.7.
@@
Let ${\mathsf{a}}\in{\mathfrak{M}}_{N-1,0}$ be an equivalence between ${\mathsf{p}}_{1}$ and ${\mathsf{p}}_{2}$, i.e. $\exp({\mathsf{a}})*{\mathsf{p}}_{2}={\mathsf{p}}_{1}.$ Let $\tilde{{\mathsf{a}}}\in{\mathfrak{M}}_{N,0}$ be any lift of ${\mathsf{a}}$. Then $r_{N,N-1}(\exp(\tilde{{\mathsf{a}}})*\tilde{{\mathsf{p}}}_{2})=r_{N,N-1}(\tilde{{\mathsf{p}}}_{1})$. By Lemma B.6, $\exp(\tilde{{\mathsf{a}}})*\tilde{{\mathsf{p}}}_{2}$ is equivalent to $\tilde{{\mathsf{p}}}_{1}$, therefore $\tilde{{\mathsf{p}}}_{2}$ is equivalent to $\tilde{{\mathsf{p}}}_{1}$. ∎
Proposition B.1.
@@
The first statement is proved by induction on $N$, where the inductive step is Lemma B.7. The second statement follows by passing to the inverse limit in $N$.
Theorem B.1.
Let ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$. Then the homology class of $[{\mathsf{p}},{\mathsf{p}}]$ in $H_{\bullet}([{\mathfrak{M}},{\mathfrak{M}}])$ is independent of the choice of ${\mathsf{p}}$.
@@
Let ${\mathsf{p}},{\mathsf{p}}^{\prime}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$ and let ${\mathsf{a}}\in{\mathfrak{M}}_{0}$ be an equivalence between ${\mathsf{p}}$ and ${\mathsf{p}}^{\prime}$. Since ${\mathsf{p}},{\mathsf{p}}^{\prime}$ satisfy the inhomogeneous Maurer-Cartan equation and ${\mathsf{B}}$ is a cycle, $[{\mathsf{p}},{\mathsf{p}}]$ and $[{\mathsf{p}}^{\prime},{\mathsf{p}}^{\prime}]$ are cycles as well. From (29) we have
${\mathsf{p}}^{\prime}-{\mathsf{p}}=-\partial{\mathsf{a}}+\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}({\mathsf{p}})}{k!}-\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}(\partial{\mathsf{a}})}{(k+1)!},$
@@
${\mathsf{f}}:=\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}({\mathsf{p}})}{k!}-\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}(\partial{\mathsf{a}})}{(k+1)!}\in{\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}.$
Thus $[{\mathsf{p}},{\mathsf{p}}]-[{\mathsf{p}}^{\prime},{\mathsf{p}}^{\prime}]=2\partial({\mathsf{p}}^{\prime}-{\mathsf{p}})=2\partial{\mathsf{f}}$ is $\partial$-exact in ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$. ∎

### v3

@@ Appendix B Inhomogeneous Maurer-Cartan equation
$\partial{\mathsf{q}}=-\frac{1}{2}[{\mathsf{p}}_{2},{\mathsf{p}}_{2}]+\frac{1}{2}[{\mathsf{p}}_{1},{\mathsf{p}}_{1}]=-[{\mathsf{q}},{\mathsf{p}}_{1}]-\frac{1}{2}[{\mathsf{q}},{\mathsf{q}}]=0.$ (161)
The last equality is because ${\mathsf{q}}\in{\mathfrak{j}}_{N}$ is central. By acyclicity of ${\mathfrak{j}}_{N}$, we have ${\mathsf{q}}=\partial{\mathsf{a}}$ for some ${\mathsf{a}}\in{\mathfrak{j}}_{N,0}$. Then Lemma 127 implies that $\exp({\mathsf{a}})\in\exp({\mathfrak{M}}_{N,0})$ maps ${\mathsf{p}}_{2}$ to ${\mathsf{p}}_{1}$. ∎
Lemma 129.
@@
Let ${\mathsf{a}}\in{\mathfrak{M}}_{N-1,0}$ be an equivalence between ${\mathsf{p}}_{1}$ and ${\mathsf{p}}_{2}$, i.e. $\exp({\mathsf{a}})*{\mathsf{p}}_{2}={\mathsf{p}}_{1}.$ Let $\tilde{{\mathsf{a}}}\in{\mathfrak{M}}_{N,0}$ be any lift of $\mathsf{a}$}. Then $r_{N,N-1}(\exp(\tilde{{\mathsf{a}}})*\tilde{{\mathsf{p}}}_{2})=r_{N,N-1}(\tilde{{\mathsf{p}}}_{1})$. By Lemma 128, $\exp(\tilde{{\mathsf{a}}})*\tilde{{\mathsf{p}}}_{2}$ is equivalent to $\tilde{{\mathsf{p}}}_{1}$, therefore $\tilde{{\mathsf{p}}}_{2}$ is equivalent to $\tilde{{\mathsf{p}}}_{1}$. ∎
Proposition 130.
@@
The first statement is proved by induction on $N$, where the inductive step is Lemma 129.
To prove the second statement, let ${\mathsf{p}},{\mathsf{p}}^{\prime}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$ and write ${\mathsf{p}}_{N}:=r_{N}({\mathsf{p}})$, ${\mathsf{p}}^{\prime}_{N}:=r_{N}({\mathsf{p}}^{\prime})\in{\rm MC}({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$. By the first part, for each $N$ there exists $g_{N}\in\exp({\mathfrak{M}}_{N,0})$ with $g_{N}*{\mathsf{p}}^{\prime}_{N}={\mathsf{p}}_{N}$.
We may choose $g_{N}$ compatibly: pick $g_{1}$ arbitrarily, and assume $g_{N-1}$ chosen. Choose any lift $\tilde{g}_{N}\in\exp({\mathfrak{M}}_{N,0})$ with $r_{N,N-1}(\tilde{g}_{N})=g_{N-1}$, and set $\tilde{\mathsf{q}}_{N}:=\tilde{g}_{N}*{\mathsf{p}}^{\prime}_{N}$. Then $r_{N,N-1}(\tilde{\mathsf{q}}_{N})=g_{N-1}*{\mathsf{p}}^{\prime}_{N-1}={\mathsf{p}}_{N-1}=r_{N,N-1}({\mathsf{p}}_{N})$, so Lemma 128 yields $h_{N}\in\exp(\ker(r_{N,N-1})_{0})$ with $h_{N}*\tilde{\mathsf{q}}_{N}={\mathsf{p}}_{N}$. Putting $g_{N}:=h_{N}\tilde{g}_{N}$ gives $r_{N,N-1}(g_{N})=g_{N-1}$ and $g_{N}*{\mathsf{p}}^{\prime}_{N}={\mathsf{p}}_{N}$.
[…]
Theorem 131.
Let ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$. Then the homology class of ${\mathsf{p}},{\mathsf{p}}]$ in $H_{\bullet}({\overline{[{\mathfrak{M}},{\mathfrak{M}}]}})$ is independent of the choice of $\mathsf{p}$}.
Proof.
Let ${\mathsf{p}},{\mathsf{p}}^{\prime}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$ and let ${\mathsf{a}}\in{\mathfrak{M}}_{0}$ be an equivalence between $\mathsf{p}$} and $\mathsf{p}$$^{\prime}$. Since ${\mathsf{p}},{\mathsf{p}}^{\prime}$ satisfy the inhomogeneous Maurer-Cartan equation and $\mathsf{B}$} is a cycle, ${\mathsf{p}},{\mathsf{p}}]$ and ${\mathsf{p}}^{\prime},{\mathsf{p}}^{\prime}]$ are cycles as well. From (158) we have
${\mathsf{p}}^{\prime}-{\mathsf{p}}=-\partial{\mathsf{a}}+{\mathsf{f}},$ (162)
@@
${\mathsf{f}}:=\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}({\mathsf{p}})}{k!}-\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}(\partial{\mathsf{a}})}{(k+1)!}\in{\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}.$ (163)
Thus ${\mathsf{p}},{\mathsf{p}}]-[{\mathsf{p}}^{\prime},{\mathsf{p}}^{\prime}]=2\partial({\mathsf{p}}^{\prime}-{\mathsf{p}})=2\partial{\mathsf{f}}$ is $\partial$-exact in $\overline{[{\mathfrak{M}},{\mathfrak{M}}]}$}. ∎